# EnScale: Temporally-consistent multivariate generative downscaling via proper scoring rules

### v2

@@ 5.4 Evaluation of performance for individual variables
5.4.1 Overall performance with scoring rules
@@
NN-det 4.5 ± 0.08 245.5 ± 1.8 245.4 0.0 1.54 ± 0.01 1.51 ± 0.01 0.9 ± 0.01
EasyUQ 6.5 ± 0.08 182.3 ± 1.47 310.1 255.9 1.22 ± 0.01 2.25 ± 0.02 6.2 ± 0.02
Analogues 7.5 ± 0.12 241.2 ± 2.7 410.6 338.1 1.68 ± 0.01 1.63 ± 0.01 1.3 ± 0.01
GAN 9.2 ± 0.09 223.4 ± 1.44 376.1 305.4 1.48 ± 0.01 1.41 ± 0.01 2.4 ± 0.02
CorrDiff 6.5 ± 0.1 212.5 ± 2.21 396.6 369.0 1.09 ± 0.01 1.06 ± 0.01 1.0 ± 0.01
EnScale 4.6 ± 0.07 168.7 ± 1.59 356.4 375.2 1.20 ± 0.01 1.17 ± 0.01 1.4 ± 0.01
EnScale-t 5.2 ± 0.09 180.2 ± 1.84 367.1 373.9 1.26 ± 0.01 1.23 ± 0.01 1.4 ± 0.01
@@
NN-det 30.9 ± 0.6 610.5 ± 5.72 610.0 0.0 2.33 ± 0.02 6.22 ± 0.06 10.1 ± 0.1
EasyUQ 49.5 ± 0.72 446.7 ± 4.48 817.2 740.6 1.94 ± 0.02 11.37 ± 0.07 9.6 ± 0.05
Analogues 39.6 ± 0.67 534.2 ± 5.96 892.3 715.9 2.35 ± 0.02 5.93 ± 0.05 5.2 ± 0.07
GAN 45.8 ± 0.9 460.5 ± 5.06 755.8 591.8 1.94 ± 0.02 5.16 ± 0.05 8.8 ± 0.07
CorrDiff 47.9 ± 0.78 414.7 ± 4.1 792.5 755.4 1.87 ± 0.02 4.68 ± 0.04 3.6 ± 0.05
EnScale 32.0 ± 0.61 434.5 ± 4.6 825.3 780.9 1.89 ± 0.02 4.85 ± 0.05 5.2 ± 0.06
EnScale-t 31.7 ± 0.62 433.7 ± 4.34 824.4 782.2 1.90 ± 0.02 4.86 ± 0.04 5.0 ± 0.06
@@
NN-det 2.8 ± 0.03 204.8 ± 1.11 204.8 0.0 1.18 ± 0.01 1.50 ± 0.01 2.1 ± 0.02
EasyUQ 4.4 ± 0.03 147.3 ± 0.8 262.8 231.1 0.92 ± 0.01 2.25 ± 0.01 9.0 ± 0.02
Analogues 4.0 ± 0.04 183.0 ± 1.53 308.9 252.0 1.14 ± 0.01 1.41 ± 0.01 2.0 ± 0.02
GAN 4.9 ± 0.04 155.8 ± 0.88 275.4 239.3 0.97 ± 0.0 1.19 ± 0.01 3.7 ± 0.03
CorrDiff 4.0 ± 0.03 141.2 ± 0.86 276.5 270.8 0.85 ± 0.0 1.04 ± 0.01 1.5 ± 0.01
EnScale 2.8 ± 0.03 138.4 ± 0.92 273.9 271.1 0.89 ± 0.0 1.09 ± 0.01 1.8 ± 0.02
EnScale-t 2.9 ± 0.03 141.2 ± 0.9 274.3 266.3 0.90 ± 0.0 1.10 ± 0.01 1.9 ± 0.02
@@
NN-det 2034.3 ± 33.68 5040.2 ± 45.05 5039.9 0.0 31.06 ± 0.26 26.14 ± 0.23 3.8 ± 0.03
EasyUQ 3359.0 ± 39.28 3614.8 ± 30.22 6765.7 6296.6 24.28 ± 0.2 47.49 ± 0.3 10.3 ± 0.03
Analogues 2628.3 ± 42.57 4360.8 ± 47.03 7303.7 5897.6 28.69 ± 0.24 20.95 ± 0.18 1.8 ± 0.02
GAN 3052.0 ± 42.32 3640.3 ± 30.95 6380.9 5470.7 23.98 ± 0.19 17.77 ± 0.15 3.9 ± 0.03
CorrDiff 3638.9 ± 57.86 3181.2 ± 27.2 6473.6 6576.6 22.84 ± 0.2 15.98 ± 0.15 1.5 ± 0.01
EnScale 2060.0 ± 32.71 3478.2 ± 34.59 7142.6 7331.4 24.11 ± 0.19 21.56 ± 0.15 2.9 ± 0.02
EnScale-t 2084.1 ± 29.71 3495.7 ± 34.16 7158.7 7331.0 24.07 ± 0.21 21.53 ± 0.15 2.8 ± 0.02
@@
a Metrics are first averaged across all days. Afterwards both metrics and standard errors are averaged across all GCM-RCM pairs to evaluate the overall performance on the entire dataset. Bold and underlined values highlight the best and second-best results, respectively.
@@
Moving from visual evaluation to more quantitative results, Tab. 3 demonstrates the performance of EnScale in comparison to the various benchmarks explained in Sec. 4.1. To assess skill overall and in each location, we show MSE, energy scores (ES), and CRPS (see Sec. 4.2.1) separately for each variable. In general, we find that EnScale reaches good results compared to the benchmarks, often achieving the best or second-best results. EnScale-t shows similar results to EnScale, but usually performs slightly worse, since these metrics only evaluate the distribution for each day independently of the previous days. A dedicated evaluation of the temporal structure follows in Sec. 5.4.5. In the following, we analyze individual metrics in more detail.
@@
The deterministic neural network excels regarding MSE for all variables. This is to be expected, as this approach explicitly minimizes the MSE during training. EnScale and EnScale-t perform only slightly worse. For all stochastic approaches (i.e., EasyUQ, analogues, GAN, CorrDiff, and EnScale), we approximate the conditional expectation $E^{\iota}[Y|X]$ as the average over nine samples corresponding to the particular GCM input $X$. With more samples per day, the estimate of the conditional mean improves, and lower scores are likely attainable, but the qualitative picture should persist.
@@
When considering the energy score averaged over multiple GCM inputs $X$, CorrDiff or EnScale perform best, with the best method depending on the variable.
@@
The energy score allows us to investigate the trade-off between prediction error ($\text{ES}_{\text{pred}}$) and variability ($\text{ES}_{\text{var}}$) quantitatively. If a method learned the correct RCM conditional distribution $Y|X$, these two scores should be equal ($\text{ES}_{\text{pred}}=\text{ES}_{\text{var}}$, see Sec. 2.5.3). For EnScale, we slightly underestimate variability for temperature and surface wind ($\text{ES}_{\text{pred}}>\text{ES}_{\text{var}}$), but overestimate variability for precipitation and solar radiation ($\text{ES}_{\text{pred}}<\text{ES}_{\text{var}}$). The latter could be a result of suboptimal convergence, whereas underestimation of variability can stem from mild overfitting. If $\text{ES}_{\text{pred}}=\text{ES}_{\text{var}}$ on the training data, $\text{ES}_{\text{pred}}$ usually increases on the test data, whereas $\text{ES}_{\text{var}}$ stays nearly constant, leading to $\text{ES}_{\text{pred}}>\text{ES}_{\text{var}}$ in test case.
@@
We find that some other benchmarks (analogues, GAN) often severely underestimate the variability compared to the prediction error ($\text{ES}_{\text{pred}}>\text{ES}_{\text{var}}$). For applications, this would imply overconfidence, i.e., a lack of spread in the generated samples, which from a risk perspective is an undesirable property of a prediction system. For GANs, for example, the lack of variance is a well-known problem [Salimans2016ImprovedGANs]. For analogues, the relationship between $\text{ES}_{\text{pred}}$ and $\text{ES}_{\text{var}}$ depends on the number of near neighbors chosen, as alluded to in Sec. 4.1. In our setting, $k=5$ neighbors still underestimate $\text{ES}_{\text{var}}$.
@@
Furthermore, we assess distributions in each location separately with the CRPS. CorrDiff reaches best scores for all variables, closely followed by EnScale. Comparison with the EasyUQ benchmark serves as another sanity check: EasyUQ is expected to perform well as it explicitly aims to best capture the pointwise CRPS values. EnScale reaches slightly better scores than the EasyUQ benchmark, showing that we successfully learned the distributions in each location despite the high-dimensional target.
@@
Scores can differ considerably between different pairs. For example, energy scores for temperature vary between 131 and 193, and those for precipitation between 359 and 562. This indicates that various RCMs disagree on the underlying variability, or possibly are not equally easy to emulate with the ML model. All above tables only consider the interpolation test set, i.e., the held-out years 2030-2039. We provide scores on the extrapolation test set (2090-2099) (see Sec. A.4). Scores are very similar and the conclusions are consistent.
@@ 5.4.2 Spatial structure
One way to assess the spatial structure is by pooling pixels spatially and computing the CPRS on these pooled results. Tab. 3 shows the results where we applied max pooling across patches of size $10\times 10$ before calculating CRPS. These findings remain consistent for other kernel sizes for pooling. The benefit of our spatially multivariate approach becomes apparent in these results. The EasyUQ benchmark, in contrast to good pointwise CRPS scores, fails to capture correct spatial structures and leads to large values for this metric. CorrDiff, EnScale and EnScale-t again show good performance and are superior to the other benchmarks.
In addition, we compute power spectral densities (PSDs) for all days and generated samples in the test set and report radially-averaged log spectral distance (RALSD, see Sec. 4.2.2). CorrDiff reaches best results, except for temperature, where NN-det has lowest RALSD. EnScale or analogues are second-best. Performance with RALSD depends on both the spatial quality as well as the similarity of samples and truth. Alternatively, we average power spectra across multiple days before calculating log spectral distances (RALSD-avg, see Sec. 4.2.2), results are shown in Tab. A.3. Here, the analogues reach lowest errors as expected because they reproduce correct spatial fields.
Plots for the PSDs in Fig. A.14 confirm that both CorrDiff and EnScale closely matches power spectra of the RCM. We find only a slight overestimation of the power for precipitation and solar radiation on small scales. Other results are as expected: EasyUQ’s have too much small-scale spatial structure, overestimating the power at small wavelengths and underestimating it at larger wavelengths. NN-det’s outputs are too smooth, generally underestimating the PSDs across wavelengths, most strikingly for precipitation and wind.
@@ 5.4.3 Calibration
We assess the calibration of the samples with rank histograms (see Sec. 4.2.3). First, we analyze calibration of spatially-aggregated quantities, the spatial mean and maximum, in Fig. 8. We find that most approaches fail to achieve good calibration for most variables, with varying degrees of violation and biases. In general, calibration for the spatial mean is often better than that for the spatial maximum.
Considering the spatial mean, we find that EnScale only shows mild violations from calibration. The samples tend to underestimate the RCM’s values for precipitation and surface wind (higher counts for larger ranks), and are slightly overdispersed for solar radiation (dome-shape of the rank histogram). However, the benchmarks do not perform better: CorrDiff shows large deviations from calibration. We find that analogues are usually underdispersed, which is in line with the scores in Tab. 3 ($\text{ES}_{\text{var}}<\text{ES}_{\text{pred}}$), and the GAN also shows errors in dispersion or biases.
@@
We compare the average miscalibration metric (MCB) across locations (see Sec. 4.2.3), see scores in the legend of Fig. 8. EnScale and EnScale-t show errors in calibration with MCB values between 0.09 and 0.17, outperforming CorrDiff. Also analogues and GAN reach higher MCB scores than EnScale (not shown). For reference, MCB scores for the spatial mean vary between 0.08 (temperature) and 0.2 (precipitation). Locationwise calibration is generally best for temperature and wind and worst for solar radiation and precipitation.
@@
These results show room for improvement for all ML approaches and point to limitations of the energy score as a loss function. In principle, a method that reaches the minimal energy score also has calibrated samples. However, in practice, the energy score is not equally sensitive to all aspects of misspecification [Pinson2013DiscriminationScore], and in some cases, a slightly sharper yet miscalibrated forecast may achieve a better energy score than a well-calibrated but less sharp one. This finding is in line with other works in weather post-processing [Wilks2018EnforcingPostprocessing, Wessel2025EnforcingModels]. It suggests that additional measures may be needed to explicitly assess and improve calibration in generative approaches.

### v3

@@ 5.3 Evaluation of performance for individual variables
5.3.1 Overall performance for conditional distributions
@@
NN-det 4.5 ± 0.08 245.5 ± 1.85 245.4 0.0 1.54 ± 0.01 1.51 ± 0.01 0.9 ± 0.01
EasyUQ 6.5 ± 0.08 182.2 ± 1.56 310.1 255.9 1.22 ± 0.01 2.25 ± 0.02 6.2 ± 0.02
Analogues 7.5 ± 0.13 241.5 ± 2.72 410.6 338.2 1.67 ± 0.01 1.63 ± 0.01 1.3 ± 0.01
GAN 9.2 ± 0.1 223.4 ± 1.46 376.0 305.3 1.48 ± 0.01 1.41 ± 0.01 2.4 ± 0.02
CorrDiff 6.5 ± 0.1 212.4 ± 2.23 396.9 368.7 1.09 ± 0.01 1.06 ± 0.01 1.0 ± 0.01
CorrDiff-s 7.7 ± 0.14 159.4 ± 1.71 319.9 321.0 1.09 ± 0.01 1.06 ± 0.01 1.0 ± 0.01
EnScale 4.6 ± 0.08 170.5 ± 1.7 351.3 361.4 1.20 ± 0.01 1.19 ± 0.01 1.3 ± 0.01
EnScale-t 4.9 ± 0.09 174.3 ± 1.76 350.6 352.8 1.23 ± 0.01 1.20 ± 0.01 1.2 ± 0.01
@@
NN-det 30.9 ± 0.61 610.3 ± 5.83 610.1 0.0 2.34 ± 0.02 6.22 ± 0.06 10.1 ± 0.1
EasyUQ 49.5 ± 0.72 447.1 ± 4.32 816.8 740.9 1.94 ± 0.02 11.37 ± 0.07 9.6 ± 0.04
Analogues 39.5 ± 0.69 535.4 ± 6.24 892.8 717.6 2.35 ± 0.02 5.93 ± 0.05 5.2 ± 0.07
GAN 45.9 ± 0.94 460.4 ± 4.88 756.0 592.1 1.94 ± 0.02 5.16 ± 0.05 8.8 ± 0.07
CorrDiff 47.9 ± 0.77 414.7 ± 4.24 792.9 756.5 1.87 ± 0.02 4.68 ± 0.04 3.6 ± 0.05
CorrDiff-s 53.5 ± 0.96 415.4 ± 3.99 812.4 794.2 1.82 ± 0.02 4.55 ± 0.05 4.2 ± 0.06
EnScale 31.9 ± 0.56 433.0 ± 4.49 825.7 785.2 1.90 ± 0.02 4.88 ± 0.05 5.1 ± 0.06
EnScale-t 31.3 ± 0.59 434.5 ± 4.45 806.0 743.5 1.89 ± 0.02 4.83 ± 0.05 5.0 ± 0.06
@@
NN-det 2.8 ± 0.03 204.7 ± 1.07 204.8 0.0 1.18 ± 0.01 1.50 ± 0.01 2.1 ± 0.02
EasyUQ 4.4 ± 0.03 147.3 ± 0.78 262.8 231.1 0.92 ± 0.0 2.25 ± 0.01 9.0 ± 0.02
Analogues 4.0 ± 0.04 183.0 ± 1.46 309.0 251.9 1.14 ± 0.01 1.41 ± 0.01 2.0 ± 0.02
GAN 4.9 ± 0.04 155.8 ± 0.86 275.4 239.1 0.97 ± 0.0 1.19 ± 0.01 3.7 ± 0.03
CorrDiff 4.0 ± 0.04 141.2 ± 0.83 276.5 270.7 0.85 ± 0.0 1.04 ± 0.01 1.5 ± 0.01
CorrDiff-s 4.7 ± 0.05 134.2 ± 0.95 264.1 260.1 0.85 ± 0.0 1.04 ± 0.01 1.6 ± 0.02
EnScale 2.8 ± 0.03 138.7 ± 0.91 272.9 268.4 0.89 ± 0.0 1.10 ± 0.01 1.8 ± 0.02
EnScale-t 2.9 ± 0.03 139.8 ± 0.91 272.5 265.2 0.89 ± 0.0 1.10 ± 0.01 1.8 ± 0.02
@@
NN-det 2033.3 ± 31.94 5036.9 ± 43.55 5041.7 0.0 31.02 ± 0.26 26.17 ± 0.21 3.8 ± 0.03
EasyUQ 3364.8 ± 39.66 3622.1 ± 33.25 6756.6 6292.3 24.32 ± 0.22 47.43 ± 0.33 10.3 ± 0.02
Analogues 2629.0 ± 37.71 4351.3 ± 48.99 7297.6 5903.9 28.70 ± 0.23 20.91 ± 0.19 1.8 ± 0.02
GAN 3049.4 ± 44.51 3646.7 ± 31.35 6377.1 5468.2 24.02 ± 0.2 17.76 ± 0.15 3.9 ± 0.03
CorrDiff 3644.4 ± 50.49 3179.7 ± 30.93 6463.3 6570.6 22.79 ± 0.2 16.00 ± 0.14 1.5 ± 0.01
CorrDiff-s 3414.7 ± 48.79 3335.0 ± 30.28 6549.1 6430.1 21.92 ± 0.2 15.38 ± 0.16 1.5 ± 0.01
EnScale 2050.7 ± 30.38 3475.0 ± 32.31 7031.4 7104.6 24.02 ± 0.2 21.05 ± 0.14 2.9 ± 0.02
EnScale-t 2064.4 ± 31.62 3463.4 ± 31.28 7060.6 7179.1 23.84 ± 0.2 21.39 ± 0.15 2.7 ± 0.02
@@
a Metrics are first averaged across all days. Afterwards both metrics and standard errors are averaged across all GCM-RCM pairs to evaluate the overall performance on the entire dataset. Bold, underlined and italic values highlight the best, second / third-best results, respectively.
@@
Moving from visual evaluation to more quantitative results, Tab. 2 demonstrates the performance of EnScale in comparison to the various benchmarks explained in Sec. 4.1. To assess skill overall and in each location, we show MSE, energy scores (ES), and CRPS (see Sec. 4.2.1) separately for each variable. In general, we find that EnScale and EnScale-t reaches good results compared to the benchmarks. One of the two methods often is among the top three methods, and the other only slightly worse. The ordering between EnScale and EnScale-t depends on the variable. A dedicated evaluation of the temporal structure follows in Sec. 5.5. In the following, we analyze individual metrics in more detail.
@@
The deterministic neural network excels regarding MSE for all variables. This is to be expected, as this approach explicitly minimizes the MSE during training. EnScale and EnScale-t perform only slightly worse. For all stochastic approaches (i.e., EasyUQ, analogues, GAN, CorrDiff/-s, and EnScale), we approximate the conditional expectation $E^{\iota}[Y|X]$ as the average over nine samples corresponding to the particular GCM input $X$. With more samples per day, the estimate of the conditional mean improves, and lower scores are likely attainable, but the qualitative picture should persist.
@@
When considering the energy score averaged over multiple GCM inputs $X$, CorrDiff or CorrDiff-s perform best, often followed by EnScale, with the best method depending on the variable.
@@
The energy score allows us to investigate the trade-off between prediction error ($\text{ES}_{\text{pred}}$) and variability ($\text{ES}_{\text{var}}$) quantitatively. If a method learned the correct RCM conditional distribution $Y|X$, these two scores should be equal ($\text{ES}_{\text{pred}}=\text{ES}_{\text{var}}$, see Sec. C.2). For EnScale, we slightly underestimate variability for precipitation and surface wind ($\text{ES}_{\text{pred}}>\text{ES}_{\text{var}}$), but overestimate variability for temperature and solar radiation ($\text{ES}_{\text{pred}}<\text{ES}_{\text{var}}$). The latter could result from suboptimal convergence, whereas underestimation of variability can stem from mild overfitting. If $\text{ES}_{\text{pred}}=\text{ES}_{\text{var}}$ on the training data, $\text{ES}_{\text{pred}}$ usually increases on the test data, whereas $\text{ES}_{\text{var}}$ stays nearly constant, leading to $\text{ES}_{\text{pred}}>\text{ES}_{\text{var}}$ in test case.
@@
We find that some other benchmarks (analogues, GAN) often severely underestimate variability relative to the prediction error ($\text{ES}_{\text{pred}}>\text{ES}_{\text{var}}$). For applications, this would imply overconfidence, i.e., a lack of spread in the generated samples, which from a risk perspective is an undesirable property of a prediction system. For GANs, for example, the lack of variance is a well-known problem [54]. For analogues, the relationship between $\text{ES}_{\text{pred}}$ and $\text{ES}_{\text{var}}$ depends on the number of near neighbors chosen, as alluded to in Sec. 4.1. In our setting, $k=5$ neighbors still underestimate $\text{ES}_{\text{var}}$.
@@
Furthermore, we assess distributions in each location separately with the CRPS. CorrDiff or CorrDiff-s reaches best scores for all variables, closely followed by EnScale or EnScale-t. Comparison with the EasyUQ benchmark serves as another sanity check: EasyUQ is expected to perform well as it explicitly aims to best capture the pointwise CRPS values. EnScale reaches slightly better scores than the EasyUQ benchmark, showing that we successfully learned the distributions in each location despite the high-dimensional target.
@@ 5.3.2 Spatial structure
Next, we assess the spatial structure of the samples. We pool pixels spatially and compute the CPRS on these pooled results. Tab. 2 shows the results where we applied max pooling across patches of size $10\times 10$ before calculating CRPS. These findings remain consistent for other kernel sizes for pooling. The benefit of our spatially multivariate approach becomes apparent in these results. The EasyUQ benchmark, in contrast to good pointwise CRPS scores, fails to capture correct spatial structures and leads to large values for this metric. CorrDiff and CorrDiff-s shows best performance, and also EnScale and EnScale-t reach low scores and are superior to the other benchmarks.
In addition, we compute power spectral densities (PSDs) for all days and generated samples in the test set and report radially-averaged log spectral distance (RALSD, see Sec. 4.2.2). CorrDiff and CorrDiff-s reach best results, except for temperature, where NN-det has lowest RALSD. EnScale or analogues are next strongest methods. Performance with RALSD depends on both the spatial quality as well as the similarity of samples and truth. Alternatively, we average power spectra across multiple days before calculating log spectral distances (RALSD-avg, see Sec. 4.2.2), results are shown in Tab. A.1. Here, the analogues reach lowest errors as expected because they reproduce correct spatial fields.
Plots for the PSDs in Fig. A.11 confirm that both EnScale and EnScale-t closely matches power spectra of the RCM. We find only a slight overestimation of the power for precipitation and solar radiation on small scales. Other results are as consistent with previous findings: CorrDiff and CorrDiff-s model power spectra accurately, with CorrDiff-s sometimes even indistinguishable from the RCM’s PSD. EasyUQ’s have too much small-scale variance, overestimating the power at small wavelengths and underestimating it at larger wavelengths. NN-det’s outputs are too smooth, generally underestimating the PSDs across wavelengths, most strikingly for precipitation and wind.
5.3.3 Marginal distribution evaluation with histograms
Figure 7: Distributions per variable for the target RCM data as well as EnScale and EnScale-t. We pool data from all locations, all GCM-RCM pairs and all days in the test set. For our methods, data is additionally pooled across 9 samples per day.
We provide a qualitative assessment of the marginal distributions of EnScale’s samples against the true RCM data with histograms (Fig. 7). We find that EnScale and EnScale-t produce very similar marginal distributions and closely match the true RCM distribution across all variables. This highlights that our multivariate approach is able to flexibly learn distributions of different climate variables despite their vastly different characteristics and shapes. The plots also suggest that both the bulk as well as the tails of the RCM’s distributions are reproduced by our ML models.
For temperature, both methods capture the tails well, with only a small underestimation of the density in the bulk of the distribution by EnScale and EnScale-t. In case of precipitation, there is broad agreement: Both EnScale and EnScale-t are able to generate very heavy rainfall events, which are also present in the true data. EnScale-t even tends to overproduce very extreme precipitation events, as suggested by the heavy upper tail. The distribution of surface wind is reproduced accurately. For solar radiation, the general shape of the distribution is well captured by our methods, although there is a small overestimation of the density at low and high values, and a small underestimation in the middle of the range.
We present histograms for some selected benchmarks in Fig. A.12. Analogues, CorrDiff and CorrDiff-s also capture the marginal distributions accurately, with small mismatches for CorrDiff, e.g., for large precipitation or high solar radiation values. The GAN, however, shows larger errors for several variables, e.g., for the mode of the surface wind or the general shape of the solar radiation distribution.
@@ 5.3.4 Calibration
A related evaluation tool are rank histograms, which assess the calibration of the conditional distributions of the samples (see Sec. 4.2.3). First, we analyze calibration of spatially-aggregated quantities, the spatial mean and maximum, in Fig. 8. We find that most approaches fail to achieve good calibration for most variables, with varying degrees of violation and biases. In general, calibration for the spatial mean is often better than that for the spatial maximum.
Considering the spatial mean, we find that EnScale only shows mild violations from calibration. The samples tend to underestimate the RCM’s values for precipitation and surface wind (higher counts for larger ranks), and are slightly overdispersed for solar radiation (dome-shape of the rank histogram). However, the benchmarks do not perform better: CorrDiff-s preforms similarly to EnScale, but CorrDiff shows large deviations from calibration. We find that analogues are usually underdispersed, which is in line with the scores in Tab. 2 ($\text{ES}_{\text{var}}<\text{ES}_{\text{pred}}$), and the GAN also shows errors in dispersion or biases.
@@
We compare the average miscalibration metric (MCB) across locations (see Sec. 4.2.3), see scores in the legend of Fig. 8. EnScale and EnScale-t show errors in calibration with MCB values between 0.08 and 0.18, similar to CorrDiff-s and outperforming CorrDiff, analogues and GAN. For reference, MCB scores for the spatial mean vary between 0.08 (temperature) and 0.2 (precipitation). Locationwise calibration is generally best for temperature and wind and worst for solar radiation and precipitation.
@@
These results show room for improvement for all ML approaches and point to limitations of the energy score as a loss function. In principle, a method that reaches the minimal energy score also has calibrated samples. However, in practice, the energy score is not equally sensitive to all aspects of misspecification [47], and in some cases, a slightly sharper yet miscalibrated forecast may achieve a better energy score than a well-calibrated but less sharp one. This finding is in line with other works in weather post-processing [79, 76]. It suggests that additional measures may be needed to explicitly assess and improve calibration in generative approaches.